OpenStudy (ali2x2):

Which points are generated by the function y = -5x + 2? Check all that apply. (2, 8) (2, -8) (13, 3) (3, -13)

OpenStudy (omarbirjas):

Thanks

OpenStudy (ali2x2):

No prob its been 7 minutes i was giving up hopr xD

OpenStudy (omarbirjas):

Anyways, this one isnt all that hard either.

OpenStudy (ali2x2):

hope if this is answered that means i have 2 left :P

OpenStudy (ali2x2):

Explain?

OpenStudy (omarbirjas):

(x, y) thats how a point is written. So all you have to do is substitute x and y in the equation for the numbers in the coordinatets that you have then solve.

OpenStudy (omarbirjas):

ex, 8=-5(2)+2

OpenStudy (omarbirjas):

then we solve, -5*2=-10

OpenStudy (omarbirjas):

-10+2=-8

OpenStudy (omarbirjas):

-8 does not equal 8

OpenStudy (omarbirjas):

we can cross out c and d since they're not equal as well, then we;re left with b.

OpenStudy (ali2x2):

Ohhh ok so the answer is b?

OpenStudy (omarbirjas):

yup.

OpenStudy (ali2x2):

Can you answer 2 more?

OpenStudy (omarbirjas):

yup but new question boxes

OpenStudy (ali2x2):

mhm ok